OpenStudy (anonymous):

Need help with a worded question (im not very good at this) if you help will give Medal + Fan (:

OpenStudy (anonymous):

@kaos_gabz

OpenStudy (anonymous):

@Chris911

OpenStudy (anonymous):

lol i tagged u and u tagged someone its like playing tag :P

OpenStudy (anonymous):

Create a quadratic equation in standard form that can be factored. start with it in factored form then multiply out for example start with \[(x-3)(x-2)\] and multiply to get \[x^2-5x+6\] then you know it can be factored

OpenStudy (chris911):

dang it he beat me

OpenStudy (anonymous):

Write your equation in factored form. Use complete sentences to explain the benefits of writing your equation in factored form. since you started it in factored form, now you know it factors as \[(x-3)(x-2)\]

OpenStudy (anonymous):

its okay @Chris911 (: and thanks @satellite73

OpenStudy (anonymous):

you can always count on chris !

OpenStudy (chris911):

u can lol

OpenStudy (anonymous):

oh no chris can u help me? satellite logged off lol

OpenStudy (chris911):

In vertex form, a quadratic function is written as y = a(x-h)2 + k

OpenStudy (chris911):

^refer to http://www.mathopenref.com/quadvertexexplorer.html

OpenStudy (anonymous):

okay sorry back was opening the website

OpenStudy (chris911):

hmm that last one is hard to explain but refer to http://jwilson.coe.uga.edu/EMT668/EMAT6680.Folders/Barron/unit/Lesson%206/6.html

OpenStudy (anonymous):

okay i got this but im like positive its probably wrong but here it is: y=3/5x^2-0.3x+9/10

OpenStudy (chris911):

this is an excellent example as a graph quadratic funtion

OpenStudy (anonymous):

Thank youuuu (:

OpenStudy (anonymous):

now i just need to make my own graph ^.^ i like graphs

OpenStudy (chris911):

u my official fan ;) wink wink lool

OpenStudy (anonymous):

hehe (: wait so was my vertex form right?

OpenStudy (chris911):

yes ;)

OpenStudy (chris911):

if not refer to a mod or get back with me and i will help u :)

OpenStudy (anonymous):

-Write your equation in vertex form. Use complete sentences to explain the benefits of writing your equation in vertex form.

OpenStudy (anonymous):

if you are going to have to write it in vertex form, i would start with a different example i would start with \[(x-2)(x-4)=x^2-6x+8\] which is factored on the left and in standard form on the right then the vertex form is easy to find since half of \(-6\) is \(-3\) it is \[(x-3)^2+k\] and we can find \(k\) by replacing \(x\) by \(3\) in the original equation and get \[k=3^2-6\times 3+8=9-18+8=-9+8=-1\] so vertex form is \[y=(x-3)^2-1\]

OpenStudy (anonymous):

Explain how all three forms can be used together to help you graph a quadratic function. Graph your function and label the y-intercept, the x-intercepts, and the vertex. \[y=(x-2)(x-4)\] is good for finding the zeros, where the graph crosses the \(x\) axis, they are \(2\) and \(4\)

OpenStudy (anonymous):

\[y=(x-3)^2-1\] is good for finding the vertex, it is \((3,-1)\)

OpenStudy (anonymous):

and \[y=x^2-6x+8\] looks pretty, it also finds the \(y\) intercept, it is \((0,8)\)

OpenStudy (anonymous):

as usual i leave the complete sentences up to you

OpenStudy (anonymous):

thank you both! (: